OpenStudy (anonymous):

a cylinder has a volume of 1538.6 cubic feet and the height of 10m feet. find the radius

OpenStudy (anonymous):

your given height is 10 meter or 10 ft?

OpenStudy (anonymous):

sorry its feet

OpenStudy (anonymous):

\[V = \pi r^2 H \] \[1538.6 ft^3 \div (10) (\pi) = r^2\] r= \[\sqrt{(1538.6 ft^3)\div(\pi)(10)}\] r=6.9982

OpenStudy (anonymous):

can you give the answer by isolating??

OpenStudy (anonymous):

I used the formula of volume which is: V=\[(\pi)r^2h\] since volume and the height is already given.. derive.....\[r=\sqrt{V \div \pi H}\] substitute the volume and the height calculate it and you'll get 6.9982